22. For a chi-square distribution,find x2\alpha

such that

a) P(X2 > x2\alpha) = 0:99 when df = 4;

b) P(X2 > x2\alpha) = 0:025 when df = 19;

c) P(37.652 < X2 < x2\alpha) = 0:045 when df = 25.


23. Find

a) t0.025 when df = 14

b) -t0.10 when df = 10

c) t0.995 when df = 7.


24. Find

a) P(T < 2:365) when df = 7;

b) P(-1.356 < T < 2.179) when df = 12;

c) P(T > 1.318) when df = 24;

d) P(T > -2.567) when df = 17.


25. For an F-distribution, find the value of f such that

a) P(F > f) = 0:05 when df1 = 4 and df2 = 9;

b) P(F > f) = 0:05 when df1 = 9 and df2 = 4;

c) P(F < f) = 0:95 when df1 = 5 and df2 = 8;

d) P(f1 < F < f2) = 0:90 when df1 = 3 and df2 = 9.


1
Expert's answer
2022-02-04T14:18:23-0500

22

"a)"

"P(\chi^2 > \chi^2_\alpha) = 0.99" when "df=4"

Here we look at "\alpha = 0.99" and "df= 4" to find from "\chi^2" tables.,

"P(X^2 > 0.297) = 0.99"

"χ^
2_{\alpha} = \chi^
2_{
0.99 }= 0.297" for "df = 4"


"b)"

"P(X^2 \gt \chi^2_{\alpha}) = 0.025" when df = 19

Here we look at "\alpha =0.025" and "df=19" to find,

"P(X^2 \gt 32.852) = 0.025"

So,

"\chi^2_{\alpha}=\chi^2_{0.025}=32.852"


"c)"

"P(37.652 \lt \chi^2 \lt \chi^2_{\alpha}) = 0.045"  when df = 25

Here we rewrite the inequality in a way that we can use the table. We Note that

"P(\chi^ 2_a \lt \chi^2 \lt \chi^2_b ) = P(\chi^ 2 \gt \chi^2_a ) − P(\chi^ 2 \gt \chi^2 _b )."

So,

"P(37.652 \lt \chi^2 \lt \chi^2_{\alpha} ) = P(\chi^ 2 \gt 37.652) − P(\chi^ 2 \gt \chi^2_{\alpha} ) = 0.045."

 First we look at the table for df= 25 and identify the "\chi^ 2_ {\alpha}"  value for 37.652 and find,

"\chi^
2_{
0.05} = 37.652"  for df = 25.

Thus,

"0.05 − P(\chi^
2 \gt \chi^2_{\alpha}) = 0.045"

or

"P(\chi^
2 \gt \chi^2_{\alpha}) = 0.005."

Now we use the table for "\alpha= 0.005" and "df = 25" to find,

 "\chi^2_{ 0.005} = 46.928" for "df = 25."


23

"a)"

"t_{0.025 }" when "df = 14"

From the t distribution table values, the t-value with "df=14" leaving an area of 0.025 to the right is 2.145.

Therefore,

"t_{{0.025,14}}=2.145"


"b)"

"-t_{0.10}" when "df = 10"

We apply the property, "-t_{\alpha}=t_{1-\alpha}".

For "-t_{0.10,10}"  we can write, "-t_{0.10,10}=t_{1-0.10,10}=t_{0.90,10}"

From tables, we have that,

"t_{0.90,10}=-1.372184" 


"c)"

"t_{0.995}" when "df = 7"

The value required here is the t distribution table value with "df = 7" that leaves an area of 0.995 to the right. From the tables, "t_{0.995,7}= -3.499483"



24

"a)"

"P(T \lt 2.365)" when "df=7"

To use the table, we must transform the inequality to,

"P(T \lt 2.365) = 1 − P(T \gt 2.365)."

Using the table for "df = 7" we see,

"t_{0.025} = 2.365," for "df= 7."

Thus,

"P(T \lt 2.365) = 1 − 0.025 = 0.975."


"b)"

"P(-1.356 \lt T \lt 2.179)" when "df = 12"

Here we use symmetry for the fact that,

"P(T \lt −t_{\alpha}) = P(T \gt t_{\alpha})."

Thus, we can write,

"P(−1.356 \lt T \lt 2.179) = 1−P(−1.356 \gt T)−P(T \gt 2.179) = 1−P(T \gt 1.356)−P(T \gt 2.179)."

Looking up these values in the table for "df = 12" yields,

"t_{0.10} = 1.356", and "t_{0.025} = 2.179" for "df = 12"

Therefore,

"P(−1.356 \lt T \lt 2.179) = 1 − 0.10 − 0.025 = 0.875."


"c)"

"P(T \gt 1.318)" when "df = 24."

Using the table directly for "df = 24" yields,

"t_{0.10} = 1.318"

Thus,

"P(T \gt 1.318) = 0.10."


"d)"

"P(T \gt -2.567)" when "df = 17"

Again, so we can use the table we write,

"P(T \gt −2.567) = 1 − P(T \lt −2.567)."

Then, symmetry gives,

"P(T \gt −2.567) = 1 − P(T \lt −2.567) = 1 − P(T \gt 2.567)"

The table gives,

"t_{0.01} = 2.567" for "df = 17"

Therefore, "P(T\gt-2.567)=1-0.01=0.99"


25

"a)"

 "P(F \gt f) = 0.05" when "df_1 = 4" and "df2 = 9" ;

Here, we look at the F distribution table value with numerator degrees of freedom equal to 4 and denominator degrees of freedom equal to 9 that leaves an area of 0.05 to the right given as,

"f_{0.05,4,9}=3.86"

So,

"P(F \gt 3.86) = 0.05" when "df_1=4" and "df_2=9"


"b)"

"P(F \gt f) = 0.05" when "df_1 = 9" and "df_2 = 4" ;

Here, we look at the F distribution table value with numerator degrees of freedom equal to 9 and denominator degrees of freedom equal to 4 that leaves an area of 0.05 to the right given as,

"f_{0.05,9,4}=6.00"

Therefore,

"P(F\gt6.00)=0.05" when "df_1=9" and "df_2=4"


"c)"

"P(F \lt f) = 0.95" when "df_1 = 5" and "df_2 = 8";

Here, "f" is a value with numerator degrees of freedom equal to 5 and denominator degrees of freedom equal to 8 that leaves an area of 0.95 to the left of the F distribution. So, this value is same as the value with numerator degrees of freedom equal to 5 and denominator degrees of freedom equal to 8 that leaves an area of 0.05 to the right of the F distribution. So, "f_{0.05,5,8}=3.69".

Therefore,

"P(F\lt 3.69)=0.95" when "df_1=5" and "df_2=8"


"d)"

"P(f_1 \lt F \lt f_2) = 0.90" when "df_1 = 3" and "df_2 = 9"

We rewrite this as, "P(f_1 \lt F \lt f_2)=P(F\gt f_1)-P(F\gt f_2)" where, the value "f_1" is the table value with "df_1=3" and "df_2=9" that leaves an area of 0.05 to the right or an area of 0.95 to the left, that is, "f_{0.95,3,9}" and "f_2" is the table value with "df_1=3" and "df_2=9" that leaves an area of 0.05 to the right, that is, "f_{0.05,3,9}".

So,

"f_2=f_{0.05,3,9}=3.86" and "f_1=f_{0.95,3,9}={1\over f_{0.05,9,3}}={1\over 8.81}=0.11"

Therefore,

"P(0.11 \lt F \lt 3.86) = 0.90" when "df_1=3" and "df_2=9."